In the linear consumption function cons = A+ Ainc The (estimated) marginal propensity to consume (MPC) out of income is simply the slope, P, while the average propensity to consume (APC) is cons /inc /inc + β. Using observations for 100 families on annual income and consumption (both is obtained: cons =-124.84+0.853 inc n=100, R2=0.692 a. (5 points) Interpret the intercept in this equation, and comment on its sign and magnitude? b. (5 points) What is the predicted consumption when family income is $30,000? c. (5 points) With inc on the x-axis, draw a graph of the estimated MPC and APC
Answer #1

(a) Intercept in the estimated equation is the value of consumption when income is zero. In the equation, this value is equal to (-124.84), which means that consumption is negative when income is zero. A negative consumption signifies savings.

(b) When Income is $30,000,

Estimated consumption ($) = - 124.84 + (0.853 x 30,000) = - 124.84 + 25,590 = 25,465.16

(c) MPC equals the coefficient of Income in the regression equation. Therefore,

MPC = 0.853

APC = Consumption / Income, therefore APC rises with increase in income. MPC and APC are graphed below.

АРС., NPI APC. МРС 0.8S3 lncome